# Robust Learning in Bayesian Parallel Branching Graph Neural Networks: The Narrow Width Limit

### v1

@@ Abstract
The infinite width limit of random neural networks is known to result in Neural Networks as Gaussian Process (NNGP) (Lee et al. (2018)), characterized by task-independent kernels. It is widely accepted that larger network widths contribute to improved generalization (Park et al. (2019)). However, this work challenges this notion by investigating the narrow width limit of the Bayesian Parallel Branching Graph Neural Network (BPB-GNN), an architecture that resembles residual networks. We demonstrate that when the width of a BPB-GNN is significantly smaller compared to the number of training examples, each branch exhibits more robust learning due to a symmetry breaking of branches in kernel renormalization. Surprisingly, the performance of a BPB-GNN in the narrow width limit is generally superior or comparable to that achieved in the wide width limit in bias-limited scenarios. Furthermore, the readout norms of each branch in the narrow width limit are mostly independent of the architectural hyperparameters but generally reflective of the nature of the data. Our results characterize a newly defined narrow-width regime for parallel branching networks in general.
@@ 1 Introduction
The study of neural network architectures has seen substantial growth, particularly in understanding how network width impacts learning and generalization. It is generally believed that wider networks generally perform better (Allen-Zhu et al. (2019); Jacot et al. (2018); Gao et al. (2024)). However, this work challenges the prevailing assumption by exploring the narrow width limit of Bayesian Parallel Branching Graph Neural Networks (BPB-GNNs), an architecture inspired by residual GCN networks (Chen et al. (2020a, 2022)). We show theoretically and empirically that narrow-width networks can perform better than their wider counterparts due to a symmetry-breaking effect in kernel renormalization, in bias-limited scenarios. This paper presents a detailed theoretical analysis of BPB-GNNs in the narrow-width regime, highlighting realistic conditions under which these networks demonstrate robust learning and comparable generalization.
@@
    We introduce a novel yet simple GCN architecture with parallel independent branches, and derive the exact generalization error for node regression in the statistical limit as the sample size $P\to\infty$ and network width $N\to\infty$, with their ratio a finite number $\alpha=P/N$, in the over-parametrized regime.
@@
    We show that in the Bayesian setting, the bias will decrease and saturate at narrow hidden layer width, a surprising phenomenon due to kernel renormalization. We demonstrate that this can be understood as a robust learning effect of each branch in the student-teacher task, where each student branch is learning the teacher’s branch.
@@
    We demonstrate this narrow-width limit in real-world dataset Cora and understand each branch’s importance as a nature of the dataset.
@@ 2 Related works
Our work follows a long tradition of mathematical analysis of infinitely-wide neural networks (Jacot et al. (2018); Lee et al. (2018); Bahri et al. (2024)), resulting in NTK or NNGP kernels. Recently, such analysis has been extended to structured neural networks including GNNs (Du et al. (2019); Walker and Glocker (2019)). However, they do not provide analysis of feature learning where the kernel is dependent on the tasks.
@@
There has been progress on understanding simple MLPs in the feature-learning regime as the kernel shape changes with the task or time (Li and Sompolinsky (2021); Atanasov et al. (2021); Avidan et al. (2023); Wang and Jacot (2023)). We develop such understanding in Graph-based networks.
@@
There is a long line of works that analyze theoretically the expressiveness (Xu et al. (2018); Geerts and Reutter (2022)) and generalization performance (Tang and Liu (2023); Garg et al. (2020); Aminian et al. (2024)) of GNN. However, it is challenging to calculate the generalization error dependence on tasks. To our knowledge, our work is first to provide a tight bound of the generalization error for GNN with residual-like structures. The closest architecture to our linear BPB-GCN is the linearly decoupled GCN proposed by Cong et al. (2021); however, the readout vector in their case has weight-sharing, which will not result in kernel renormalization for different branches.
@@ 3 BPB-GNN
We are motivated to study the parallel branching networks as they resemble residual blocks in commonly used architectures and tractable to study analytically with our Bayesian framework. Given graph $G=(A,X)$, where $A$ is the adjacency matrix and $X$ the node feature matrix, the final readout for node $\mu$ is a scalar $f^{\mu}(G;\Theta)$ which depends on the graph and network parameters $\Theta$.
@@ 3.1 Parallel branching GNN architecture
The parallel branching GNN is an ensemble of GNN branches, where each branch operates independently with no weight sharing. In this work, we focus on the simple setup of branches made of linear GCN with one hidden layer, but with different number of convolutions $A$ [^1] In this paper, the convolution operation is normalized as $A=D^{-1/2}(\hat{A}+I)D^{-1/2}$, where $\hat{A}$ is the original Adjacency matrix and $D$ is the degree matrix. We also use feature standardization after convolution for each branch to normalize the input. on the input node features (Figure 1(a)). In this way, the parallel branching GNN is analogous to GCN with residual connections, for which the final node readout can also be thought of as an ensemble of convolution layers (Veit et al. (2016)). Concretely, the overall readout $f^{\mu}(G;\Theta)$ for node $\mu$ is a sum of the branch readouts
@@
Note that when $L=2$, the BPB-GNN reduces exactly to a 2-layer residual GCN (Chen et al. (2020b)).
(a) The parallel branching GNN architecture, with 2 branches.
(b) Student and teacher readout norms squared for wide and narrow student BPB-GNN networks.
Figure 1: Overview of the main message of the paper: narrow BPB-GNN learns robust representations for each branch at narrow width.
@@ 3.2 Bayesian node regression
We consider a Bayesian semi-supervised node regression problem, for which the posterior probability for the weight parameters is given by
$P(\Theta)=\frac{1}{Z}e^{-E(\Theta;G,Y)/T}=\frac{1}{Z}\exp({-\frac{1}{2T}\sum_{\mu=1}^{P}(f^{\mu}(G,\Theta)-y^{\mu})^{2}-\frac{1}{2\sigma^{2}}\Theta^{T}\Theta}),$ (4)
@@
We are interested in understanding the weight and predictor statistics of each branch and how they contribute to the overall generalization performance of the network. In the following theoretical derivations, our working regime is in the overparametrizing high dimensional limit (Li and Sompolinsky (2021); Montanari and Subag (2023); Bordelon and Pehlevan (2022); Howard et al. (2024)): $P,N,N_{0}\to\infty$, $\frac{P}{N}=\alpha$ finite, and the capacity $\alpha_{0}=\frac{P}{LN_{0}}$<1. As we will show later, this limit is practically true even with $P,N$ not so large (our smallest $N$ is 4). We will also use near 0 temperature in which case the training error will be near 0 and the prior $L_{2}$ regularization has an inductive bias on the solution space that will influence the generalization properties.
@@ 3.3 Kernel renormalization and order parameters
where the effective Hamiltonian $H(W)$ in terms of the hidden layer weights for all branches is
$H(W)=\frac{1}{2\sigma^{2}}\sum_{l=0}^{L-1}\mathrm{Tr}W_{l}^{T}W_{l}+\frac{1}{2}Y^{T}(K(W)+TI)^{-1}Y+\frac{1}{2}\log\det(K(W)+TI),$ (6)
@@
$K(W)=\frac{1}{L}\sum_{l}\frac{\sigma^{2}}{N}(H_{l}(W_{l})H_{l}(W_{l})^{T})|_{P}$ (7)
@@
$S(u)=-\sum_{l}\frac{N}{2}\log u_{l}+\sum_{l}\frac{N}{2\sigma^{2}}u_{l}$ (9)
@@
$E(u)=\frac{1}{2}Y^{T}(\sum_{l}\frac{1}{L}u_{l}K_{l}+TI)^{-1}Y+\frac{1}{2}\log\det(\sum_{l}\frac{1}{L}u_{l}K_{l}+TI),$ (10)
where $K_{l}=\frac{\sigma^{2}}{N_{0}}[A^{l}XX^{T}A^{l}]|_{P}$ is the ($P\times P$) input node feature kernel.
@@
where $u_{l}$’s are order parameters which is the minimum of the effective Hamiltonian Eq. 8 by saddle point methods.
@@
Observe that as $\alpha=P/N\to 0$, the entropic term dominates, and thus $u_{l}=\sigma^{2}$ for all branches. This is the usual Gaussian Process (GP) limit, or the infinite-width limit. In this case, the kernel $K_{l}$ for each branch will not be changed by the training data and each branch has the same contribution in terms of its strength.
However, when $\alpha=P/N$ is large, there is going to be a correction to the GP prediction and $u_{l}$’s in general are going to depend on the training data; we therefore have feature learning in each branch with kernel renormalization. It turns out that $u_{l}$ is exactly the statistical average of readout norm squared (Appendix A.3)
@@
for each branch $l$; therefore, branches in general become more and more different as the hidden layer width $N$ gets smaller. We call this phenomenon symmetry breaking, which is discussed in Section 4.2.
@@ 3.4 Predictor statistics and generalization
$\langle y^{\nu}(G)\rangle=\sum_{l}\frac{1}{L}u_{l}k_{l,\nu}(K+TI)^{-1}Y$ (13)
$\langle\delta y_{\nu}(G)^{2}\rangle=K_{\nu,\nu}-\sum_{l^{\prime}}\frac{1}{L}u_{l^{\prime}}k_{l^{\prime},\nu}^{T}(K+TI)^{-1}\sum_{l^{\prime\prime}}\frac{1}{L}u_{l^{\prime\prime}}k_{l^{\prime\prime},\nu}$ (14)
@@
$K^{\nu,\nu}=\frac{\sigma^{2}}{N_{0}}\sum_{l}\frac{1}{L}u_{l}[A^{l}XX^{T}A^{l}]|_{(\nu,\nu)}$ (16)
@@ 4 The narrow width limit
As we discussed briefly in 3.3, the kernel becomes highly renormalized at narrow width. In fact, in the other extreme scenario when $N/P\to 0$, the energetic term in the Hamiltonian completely dominates, and we would expect that the generalization performance saturates as the order parameters in the energetic terms become independent of width $N$. Therefore, just as infinitely wide networks correspond to the GP limit, we propose that there exists a narrow width limit when the network width is extremely small compared to the number of training samples.
@@ 4.1 Robust learning of branches: the equipartition conjecture
$f^{*}(G;\Theta^{*})=\sum_{l}f_{l}^{*}(G;W_{l}^{*})=\sum_{l}\frac{1}{\sqrt{NL}}\sum_{i}a_{i,l}^{*}h_{i}^{l}(G;W_{l}^{*}).$ (19)
$W_{ij}^{*}\sim\mathcal{N}(0,\sigma_{t}^{2})$ and $a_{i,l}^{*}\sim\mathcal{N}(0,\beta_{l}^{2})$, where $\beta_{l}^{2}$ is the assigned variance of the readout weight for the teacher branch $l$. Now we feed the data generated from this distribution $Y$ to the student network with the same number of branches, with student variance $\sigma^{2}$.
Consider $(YY^{T})_{\mu,\nu}=\sum_{i,j,l_{1},l_{2}}\frac{1}{NL}a^{*}_{i,l_{1}}a^{*}_{j,l_{2}}h_{i}^{l_{1}}(G)h_{j}^{l_{2}}(G)$; we conjecture that this quantity concentrates at its expectation value $\mathbf{E}_{a^{*},W^{*}}(YY^{T})=\sum_{l}\beta_{l}^{2}K_{l}/L$. Given this assumption, we can calculate the order parameters in the saddle point equation 40 (A.1) at 0 temperature. Ignoring the entropic term at narrow width, $r_{l}$ (Appendix A) becomes
$r_{l}=Y^{T}K^{-1}\frac{u_{l}K_{l}}{L}K^{-1}Y\approx\mathrm{Tr}K^{-1}\frac{u_{l}K_{l}}{L}(K^{-1}\sum_{l}\beta_{l}^{2}\frac{K_{l}}{L})=\mathrm{tr}_{l}(K^{-1}\sum_{l}\beta_{l}^{2}\frac{K_{l}}{L})$ (20)
One solution that satisfies the saddle point equations is
$u_{l}\sigma^{2}=\beta_{l}^{2}\sigma_{t}^{2}.$ (21)
We call this the equipartition conjecture, as the mean readout squared and the variance (A.3) has to exactly balance each other. Furthermore, by Eq.12, the student readout norm at each branch on average learns the teacher readout norm when the variance of teacher network $\sigma_{t}^{2}=\sigma^{2}$ and $N\to\infty$, ie.
$\langle\|a_{\mathrm{student}}\|^{2}\rangle=\|a_{\mathrm{teacher}}\|^{2}$ (22)
Furthermore, writing $u_{l}K_{l}/L\approx Y_{l}Y^{T}$, at narrow width, the predictor statistics for each stream for the training data from the teacher becomes
$\langle f_{l}(X)\rangle=\frac{u_{l}K_{l}}{L}(K+TI)^{-1}Y\approx Y_{l},$ (23)
ie. not only do we recover the statistics for $a_{l}$’s, we have also recovered the feature learned by each branch.

### v2

@@ Abstract
The infinite width limit of random neural networks is known to result in Neural Networks as Gaussian Process (NNGP) (Lee et al. (2018)), characterized by task-independent kernels. It is widely accepted that larger network widths contribute to improved generalization (Park et al. (2019)). However, this work challenges this notion by investigating the narrow width limit of the Bayesian Parallel Branching Graph Neural Network (BPB-GNN), an architecture that resembles residual GCN. We demonstrate that when the width of a BPB-GNN is significantly smaller compared to the number of training examples, each branch exhibits more robust learning due to a symmetry breaking of branches in kernel renormalization. Surprisingly, the performance of a BPB-GNN in the narrow width limit is generally superior to or comparable to that achieved in the wide width limit in bias-limited scenarios. Furthermore, the readout norms of each branch in the narrow width limit are mostly independent of the architectural hyperparameters but generally reflective of the nature of the data. Our results characterize a newly defined narrow-width regime for parallel branching networks in general.
@@ 1 Introduction
The study of neural network architectures has seen substantial growth, particularly in understanding how network width impacts learning and generalization. In general, wider networks are believed to perform better (Allen-Zhu et al. (2019); Jacot et al. (2018); Gao et al. (2024)). However, this work challenges the prevailing assumption by exploring the narrow width limit of Bayesian Parallel Branching Graph Neural Networks (BPB-GNNs), an architecture inspired by residual GCN networks (Chen et al. (2020a; 2022)). We show theoretically and empirically that narrow-width networks can perform better than their wider counterparts due to a symmetry-breaking effect in kernel renormalization, in bias-limited scenarios. This paper presents a detailed theoretical analysis of BPB-GNNs in the narrow-width regime, highlighting realistic conditions under which these networks demonstrate robust learning and comparable generalization.
@@
    We introduce a novel yet simple GCN architecture with parallel independent branches, and derive the exact generalization error for node regression in the statistical limit as the sample size $P\to\infty$ and network width $N\to\infty$, with their ratio a finite number $\alpha=P/N$, in the over-parameterized regime.
@@
    We show that in the Bayesian setting the bias will decrease and saturate at a narrow hidden layer width, a surprising phenomenon due to kernel renormalization. We demonstrate that this can be understood as a robust learning effect of each branch in the student-teacher task, where each student’s branch is learning the teacher’s branch.
@@
    We demonstrate this narrow-width limit in the real-world dataset Cora and understand each branch’s importance as a nature of the dataset.
@@ 2 Related works
Our work follows a long tradition of mathematical analysis of infinitely-wide neural networks (Neal (2012); Jacot et al. (2018); Lee et al. (2018); Bahri et al. (2024)), resulting in NTK or NNGP kernels. Recently, such analysis has been extended to structured neural networks, including GNNs (Du et al. (2019); Walker & Glocker (2019)). However, they do not provide an analysis of feature learning in which the kernel depends on the tasks.
@@
There has been progress in understanding simple MLPs in the feature-learning regime as the shape of the kernel changes with task or time (Li & Sompolinsky (2021); Atanasov et al. (2021); Avidan et al. (2023); Wang & Jacot (2023)). We develop such understanding in graph-based networks.
@@
There is a long line of works that theoretically analyze the expressiveness (Xu et al. (2018); Geerts & Reutter (2022)) and generalization performance (Tang & Liu (2023); Garg et al. (2020); Aminian et al. (2024)) of GNN. However, it is challenging to calculate the dependence of generalization errors on tasks. To our knowledge, our work is first to provide a tight bound of the generalization error for GNN with residual-like structures. The architecture closest to our linear BPB-GCN is the linearly decoupled GCN proposed by Cong et al. (2021); however, the overall readout vector is shared for all branches, which will not result in kernel renormalization for different branches.
@@ 3.1 Parallel branching GNN architecture
We are motivated to study the parallel branching networks as they resemble residual blocks in commonly used architectures (Kipf & Welling (2016); He et al. (2016); Chen et al. (2020b)) and are easy to study analytically with our Bayesian framework. Given graph $G=(A,X)$, where $A$ is the adjacency matrix and $X$ the node feature matrix, the final readout for node $\mu$ is a scalar $f^{\mu}(G;\Theta)$ that depends on the graph and network parameters $\Theta$. The parallel branching GNN is an ensemble of GNN branches, where each branch operates independently with no weight sharing. In this work, we focus on the simple setup of branches made of linear GCN with one hidden layer, but with different number of convolutions $A$ [^1] In this paper, the convolution operation is normalized as $A=D^{-1/2}(\hat{A}+I)D^{-1/2}$, where $\hat{A}$ is the original Adjacency matrix and $D$ is the degree matrix. We also use feature standardization after convolution for each branch to normalize the input on the input node features (Figure 1(a)). In this way, parallel branching GNN is analogous to GCN with residual connections, for which the final node readout can also be thought of as an ensemble of convolution layers (Veit et al. (2016)). Concretely, the overall readout $f^{\mu}(G;\Theta)$ for node $\mu$ is a sum of $L$ branch readouts $f_{l}^{\mu}$:
@@
Note that when $L=2$, the network reduces exactly to a 2-layer residual GCN (Chen et al. (2020c)). Here $N_{0}$ is the input dimension, $N$ the width of the hidden layer, $W^{(l)}$ the weight of the hidden layer for branch $l$ and $a^{(l)}$ the final readout weight for branch $l$. We will consider only the linear activation function for this paper and provide a summary of notations in Appendix A.1.
Refer to caption Refer to caption
(a) (b)
Figure 1: Overview of the main takeaway: BPB-GNN learns robust representations for each branch at narrow width. (a) The parallel branching GNN architecture, with 2 branches. The independent branches have non-sharing weights and produce the final output $f$ as a sum of branch-level readouts $f_{l}$. (b) Student and teacher readout norms squared for wide and narrow student BPB-GNN networks. The student network with width $N$ is trained with the teacher network’s output. Histograms correspond to the samples from Hamiltonian Monte Carlo simulations and solid lines correspond to the order parameters calculated theoretically. $\sigma_{t}=\sigma_{w}=1$. At $N=4$, the HMC samples of branch readout norms squared (orange and red histograms) for the student network $\frac{\|a_{l}\|^{2}}{N}\sigma_{w}^{2}$ concentrate at their respective theoretical values $u_{l}\sigma_{w}^{2}$ and overlap with the teacher’s readout norms squared $\frac{\|A_{l}\|^{2}}{N}\sigma_{t}^{2}$ (orange and red dashed lines) for corresponding branches. At $N=1024$ the samples for the student network (blue and green histograms) concentrate at their respective theoretical values but remain far from the teacher’s values, instead approaching the GP limit $\sigma_{w}^{4}$ (blue dashed line).
@@ 3.2 Bayesian node regression
We consider a Bayesian semi-supervised node regression problem, for which the posterior probability distribution for the weight parameters is given by
$P(\Theta)=\frac{1}{Z}e^{-E(\Theta;G,Y)/T}=\frac{1}{Z}\exp({-\frac{1}{2T}\sum_{\mu=1}^{P}(f^{\mu}(G,\Theta)-y^{\mu})^{2}-\frac{1}{2\sigma_{w}^{2}}\Theta^{T}\Theta}),$ (4)
@@
We are interested in understanding the weight and predictor statistics of each branch and how they contribute to the overall generalization performance of the network. In the following theoretical derivations, our working regime is in the overparameterizing high-dimensional limit (Li & Sompolinsky (2021); Montanari & Subag (2023); Bordelon & Pehlevan (2022); Howard et al. (2024)): $P,N,N_{0}\to\infty$, $\frac{P}{N}=\alpha$ finite and the capacity $\alpha_{0}=\frac{P}{LN_{0}}<1$. As we will show later, this limit is practically true even with $P,N$ not so large (our smallest $N$ is 4). We will also use near-0 temperature, in which case the training error will be near 0.
@@ 3.3 Kernel renormalization and order parameters
with effective Hamiltonian $H(W)$ in terms of the hidden layer weights for all branches
$H(W)=\frac{1}{2\sigma_{w}^{2}}\sum_{l=0}^{L-1}\mathrm{Tr}W_{l}^{T}W_{l}+\frac{1}{2}Y^{T}(K(W)+TI)^{-1}Y+\frac{1}{2}\log\det(K(W)+TI),$ (6)
@@
$K(W)=\frac{1}{L}\sum_{l}\frac{\sigma_{w}^{2}}{N}(H_{l}(W_{l})H_{l}(W_{l})^{T})|_{P}$ (7)
@@
$S(u)=-\sum_{l}\frac{N}{2}\log u_{l}+\sum_{l}\frac{N}{2\sigma_{w}^{2}}u_{l}$ (9)
@@
$E(u)=\frac{N\alpha}{2P}Y^{T}(\sum_{l}\frac{1}{L}u_{l}K_{l}+TI)^{-1}Y+\frac{N\alpha}{2P}\log\det(\sum_{l}\frac{1}{L}u_{l}K_{l}+TI),$ (10)
where $K_{l}=\frac{\sigma^{2}}{N_{0}}[A^{l}XX^{T}A^{l}]|_{P}$ is the ($P\times P$) input node feature kernel for branch $l$.
@@
where $u_{l}$’s are order parameters which are the minimum of the effective Hamiltonian Eq. 8 that satisfy the saddle point equations:
$N(1-\frac{u_{l}}{\sigma_{w}^{2}})=-r_{l}+\mathrm{Tr}_{l},$ (12)
where $r_{l}=Y^{T}(K+TI)^{-1}\frac{u_{l}K_{l}}{L}(K+TI)^{-1}Y$ and $\mathrm{Tr}_{l}=\mathrm{Tr}[K^{-1}\frac{u_{l}K_{l}}{L}]$.
@@
Observe that as $\alpha=P/N\to 0$, the entropic term dominates, and thus $u_{l}=\sigma_{w}^{2}$ for all branches from Eq. 12. This is the usual Gaussian Process (GP) limit, or the infinite-width limit. In this case, the kernel $K_{l}$ for each branch is not changed by the training data, and each branch has the same contribution in terms of its strength.
However, when $\alpha=P/N$ is large, there is a correction to the GP prediction, and $u_{l}$’s in general depend on the training data; therefore, we have feature learning in each branch with kernel renormalization. It turns out that $u_{l}$ is exactly the statistical average of readout norm squared over the posterior distribution (Appendix A.4), ie.
@@
for each branch $l$. Therefore, branches in general become more and more different as the width of the hidden layer $N$ decreases. We call this phenomenon symmetry breaking, which is discussed in Section 4.2.
@@ 3.4 Predictor statistics and generalization
$\langle f^{\nu}(G)\rangle=\sum_{l}\frac{1}{L}u_{l}k_{l,\nu}(K+TI)^{-1}Y$ (14)
$\langle\delta f_{\nu}(G)^{2}\rangle=K_{\nu,\nu}-\sum_{l^{\prime}}\frac{1}{L}u_{l^{\prime}}k_{l^{\prime},\nu}^{T}(K+TI)^{-1}\sum_{l^{\prime\prime}}\frac{1}{L}u_{l^{\prime\prime}}k_{l^{\prime\prime},\nu}$ (15)
@@
$K^{\nu,\nu}=\frac{\sigma_{w}^{2}}{N_{0}}\sum_{l}\frac{1}{L}u_{l}[A^{l}XX^{T}A^{l}]|_{(\nu,\nu)}$ (17)
@@ 4 The narrow width limit
As we discussed briefly in Section 3.3, the kernel becomes highly renormalized at narrow width. In fact, in the extreme scenario as $\alpha=P/N\to\infty$, the energetic term in the Hamiltonian completely dominates, and we would expect that the generalization performance saturates as the order parameters in the energetic terms become independent of width $N$. Therefore, just as infinitely wide networks correspond to the GP limit, we propose that there exists a narrow width limit when the network width is extremely small compared to the number of training samples.
@@ 4.1 Robust learning of branches: the equipartition conjecture
$f^{*}(G;\Theta^{*})=\sum_{l}f_{l}^{*}(G;W_{l}^{*})=\sum_{l}\frac{1}{\sqrt{N_{t}L}}\sum_{i}A_{i,l}h_{i}^{l}(G;W_{l}^{*}).$ (20)
$W_{ij}^{*}\sim\mathcal{N}(0,\sigma_{t}^{2})$ and $A_{i,l}\sim\mathcal{N}(0,\beta_{l}^{2})$, where $\beta_{l}^{2}$ is the variance assigned to the readout weight for the teacher branch $l$ and $N_{t}$ is the width of the hidden layer. Similarly, the student network is given by the same architecture, with layer width $N$ and learns from $P$ node lables from the teacher $Y_{\mu=1}^{*P}=f^{*}(G,\Theta^{*})_{\mu}$ in the Bayesian regression setup of Eq. 4 with prior variance $\sigma_{w}^{2}$. We conjecture that as $\alpha=P/N\to\infty$ the posterior distribution of the student network readout vector $a_{l}$ satisfies
$\sigma_{w}^{2}\langle\|a_{l}\|^{2}\rangle=\sigma_{t}^{2}\|A_{l}\|^{2}.$ (21)
Sketch of proof:
At narrow width, the saddle equation 12 becomes $r_{l}=\mathrm{Tr}_{l}$. Now consider $(Y^{*}Y^{*T})_{\mu,\nu}=\sum_{i,j,l_{1},l_{2}}\frac{1}{N_{t}L}A_{i,l_{1}}A_{j,l_{2}}h_{i}^{l_{1},\mu}(G)h_{j}^{l_{2},\nu}(G)$; we conjecture that this quantity concentrates at its expectation value $Y^{*}Y^{*T}\approx\mathbb{E}_{a^{*},W^{*}}(YY^{T})=\sum_{l}\beta_{l}^{2}K_{l}/L$. Given this assumption, $r_{l}$ becomes
$r_{l}=Y^{*T}K^{-1}\frac{u_{l}K_{l}}{L}K^{-1}Y^{*}=\mathrm{Tr}(K^{-1}\frac{u_{l}K_{l}}{L}Y^{*}Y^{*T})\approx\mathrm{Tr}_{l}(K^{-1}\sum_{l}\beta_{l}^{2}\frac{K_{l}}{L})$ (22)
The solution that satisfies the saddle point equations is
$u_{l}\sigma_{w}^{2}=\beta_{l}^{2}\sigma_{t}^{2}.$ (23)
Therefore, by Eq.13, we proved the conjecture given $N_{t}$ is also large enough. We call this equipartition conjecture, as the mean-squared readout and the variance (A.4) have to exactly balance each other, which contribute to the energy term in the Hamiltonian. This is only a conjecture, as it relies on the concentration equality assumption made in the proof. However, as we demonstrate in the experiments, the equipartition conjecture holds empirically.
Furthermore, writing $u_{l}K_{l}/L\approx Y^{*}_{l}Y^{*T}$, at narrow width, the predictor statistics for each stream for the training data from the teacher becomes
$\langle f_{l}(X)\rangle=\frac{u_{l}K_{l}}{L}(K+TI)^{-1}Y^{*}\approx Y^{*}_{l},$ (24)
ie. not only do we recover the statistics for the teacher $A_{l}$’s, we have also recovered the feature learned by each branch.